OpenStudy (anonymous):

Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right.

OpenStudy (bahrom7893):

4x - 7 = 5

OpenStudy (bahrom7893):

Add 7 to both sides: 4x - 7 + 7 = 5 + 7 4x = 12 Divide both sides by 4: 4x/4 = 12/4 x = 3

OpenStudy (anonymous):

Thank you sssssoooooo much!!

OpenStudy (bahrom7893):

And we can check the answer easily: 4x-7=5 4*3 - 7 = 12 - 7 = 5 = 5, it works